# A quasilinear Keller-Segel model with saturated discontinuous advection

### v2

@@ 4.2. Identification of the limit of $\nabla\rho_{\varepsilon}$
almost everywhere in $[0,T]\times\mathbb{R}^{d}\setminus\{\rho=0\}$. But as $\rho_{\varepsilon}\rightarrow\rho$ strongly in $L^{2}(0,T,\mathbb{R}^{d})$, $\rho_{\varepsilon}\rightarrow\rho$ a.e. in $(0,T)\times\mathbb{R}^{d}$, and
@@
almost everywhere in $[0,T]\times\mathbb{R}^{d}\setminus\{\rho=0\}$.
@@
$\int_{Q_{R,\tau,T}^{\eta}}\left[\left(|\nabla\rho_{\varepsilon}|^{2}-\rho^{2}\chi^{2}\right)_{+}\nabla\rho_{\varepsilon}-\left(|\nabla\rho|^{2}-\rho^{2}\chi^{2}\right)_{+}\nabla\rho\right](\nabla\rho_{\varepsilon}-\nabla\rho_{\varepsilon})\,{\rm d}x\,{\rm d}t$
@@
$+\frac{1}{2}\int_{Q_{R,\tau,T}^{\eta}}\left[\left(|\nabla\rho_{\varepsilon}|^{2}-\rho^{2}\chi^{2}\right)_{+}-\left(|\nabla\rho|^{2}-\rho^{2}\chi^{2}\right)_{+}\right]\times\left(\lvert\nabla\rho_{\varepsilon}\rvert^{2}-\lvert\nabla\rho_{\varepsilon}\rvert^{2}\right)\,{\rm d}x\,{\rm d}t\rightarrow 0.$ (41b)
Since both integrals in (41a), (41b) are nonnegative, the convergence to 00 of (41a) implies that we can find a subsequence (not relabeled) such that $\nabla\rho_{\varepsilon}\rightarrow\nabla\rho$ a.e. in $Q_{R,\tau,T}^{\eta}\setminus\{\rho\chi\geq|\nabla\rho|\}$. W.l.o.g. we can choose $\eta=1/N$, $N\in\mathbb{N}$, and assume that $Q_{R,\tau,T}^{1/N_{1}}\subset Q_{R,\tau,T}^{1/N_{2}}$ for $N_{1}<N_{2}$. Since $|Q_{R,\tau,T}\setminus Q_{R,\tau,T}^{\eta}|<\eta$, a Cantor diagonal argument yields the existence of a subsequence (not relabeled) such that $\nabla\rho_{\varepsilon}\rightarrow\nabla\rho$ a.e. in $Q_{R,\tau,T}\setminus\{\rho\chi\geq|\nabla\rho|\}$. Via another Cantor diagonal argument we conclude that (up to a subsequence)
$\nabla\rho_{\varepsilon}\rightarrow\nabla\rho\leavevmode\nobreak\ \leavevmode\nobreak\ \mbox{ a.e. in }(0,T)\times\mathbb{R}^{d}\setminus\{\rho\chi\geq|\nabla\rho|\}.$ (42)
@@
$\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}\nabla\rho_{\varepsilon}\rightharpoonup\left(1-\frac{\chi\rho}{\lvert\nabla\rho\rvert}\right)_{+}\nabla\rho\quad\mbox{weakly in }L^{2}((0,T)\times\mathbb{R}^{d}\setminus\{\rho\chi\geq|\nabla\rho|\}).$
@@
$\left[\left(|\nabla\rho_{\varepsilon}|^{2}-\rho^{2}\chi^{2}\right)_{+}-\left(|\nabla\rho|^{2}-\rho^{2}\chi^{2}\right)_{+}\right]\times\left(\lvert\nabla\rho_{\varepsilon}\rvert^{2}-\lvert\nabla\rho_{\varepsilon}\rvert^{2}\right)$
@@
$\left(|\nabla\rho_{\varepsilon}|^{2}-\rho_{\varepsilon}^{2}\chi^{2}\right)_{+}\rightarrow\left(|\nabla\rho|^{2}-\rho^{2}\chi^{2}\right)_{+}\quad\mbox{a.e. in }L^{2}({(0,T)\times\mathbb{R}^{d}\cap\{\rho>0\}}).$
Furthermore, notice that
@@
a.e. in $(0,T)\times\mathbb{R}^{d}\cap\{\rho\chi\geq|\nabla\rho,|\leavevmode\nobreak\ {\rho>0}\}$. Finally, it holds
$\left|\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}\nabla\rho_{\varepsilon}\right|\leq 2\sqrt{\rho_{\varepsilon}}|\nabla\sqrt{\rho_{\varepsilon}}|\to 0=\left|\left(1-\frac{\chi\rho}{\lvert\nabla\rho\rvert}\right)_{+}\nabla\rho\right|\quad\mbox{a.e.\leavevmode\nobreak\ in }\{\rho>0\}.$
This means that
@@
and we finally get
$\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}\nabla\rho_{\varepsilon}\rightharpoonup\left(1-\frac{\chi\rho}{\lvert\nabla\rho\rvert}\right)_{+}\nabla\rho\quad\mbox{weakly in }L^{2}((0,T)\times\mathbb{R}^{d}).$
4.3. Limit $\varepsilon\rightarrow 0$ in (3.2)
As in (19), we have that
$\|\partial\rho_{\varepsilon}\|_{L^{2}(0,T;H^{-1}(\mathbb{R}^{d}))}\leq C\|\rho_{{\rm in},\varepsilon}\|_{L^{2}(\mathbb{R}^{d})},$
and, as $\rho_{{\rm in},\varepsilon}$ is uniformly bounded in $L^{2}(0,T;\mathbb{R}^{d})$, this yields
@@
On the other hand, due to what is stated in Section 4, we have that
$\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}\nabla\rho_{\varepsilon}\rightharpoonup\left(1-\frac{\chi\rho}{\lvert\nabla\rho\rvert}\right)_{+}\nabla\rho\quad\mbox{weakly in }\;\;L^{2}(0,T;\mathbb{R}^{d}).$
For the regularizing terms we have the following:
$\lim_{\varepsilon\rightarrow 0}\left|\varepsilon\int_{0}^{T}\int_{\mathbb{R}^{d}}\nabla\rho_{\varepsilon}\nabla\varphi\,{\rm d}x\,{\rm d}t+\varepsilon\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\varphi\,{\rm d}x\,{\rm d}t\right|$ $\leq\lim_{\varepsilon\rightarrow 0}\varepsilon C\|\rho_{{\rm in},\varepsilon}\|_{L^{2}(0,T;\mathbb{R}^{d})}\|\varphi\|_{L^{2}(0,T;\mathbb{R}^{d})}$
@@
Finally we have that $\rho$ is the weak solution to (10).
4.4. Obtaining the initial data $\rho_{{\rm in}}$ for $\rho$
@@
$\left\{u\in L^{2}(0,T;H^{1}(\mathbb{R}^{d}))\leavevmode\nobreak\ \leavevmode\nobreak\ :\leavevmode\nobreak\ \leavevmode\nobreak\ \partial_{t}u\in L^{2}(0,T;H^{-1}(\mathbb{R}^{d}))\right\}$
@@
[…]
Also remember that the initial condition for the approximated problem (3.2) is satisfied in the sense of (18).
Define for brevity $A_{\varepsilon}:=\varepsilon+\left(1-\frac{\chi\rho_{\varepsilon}}{|\nabla\rho_{\varepsilon}|}\right)_{+}$, $A:=\left(1-\frac{\chi\rho}{|\nabla\rho|}\right)_{+}$. Choose $\varphi\in H^{1}(\mathbb{R}^{d})$ independent of time in (3.2). It follows

### v3

@@ 4.2. Identification of the limit of the flux
$\frac{\nabla\rho_{\varepsilon}}{\rho_{\varepsilon}}{\bf 1}_{\{\rho_{\varepsilon}>0\}}\rightarrow w\quad\mbox{a.e. in }[0,T]\times\mathbb{R}^{d}\setminus\{\{\rho=0\}\cup\{\Xi=0\}\}.$ (43)
Notice in particular that the a.e. convergence of ${\bf 1}_{\{\rho_{\varepsilon}>0\}}\nabla\rho_{\varepsilon}$ in $\{\rho>0,\Xi>0\}$ follows (since $\rho_{\varepsilon}$ is a.e. convergent). Given that ${\bf 1}_{\{\rho_{\varepsilon}=0\}}\nabla\rho_{\varepsilon}=0$ a.e. in $\mathbb{R}^{d}\times(0,T)$, we infer the a.e. convergence of $\nabla\rho_{\varepsilon}$ on $\{\rho>0,\,\Xi>0\}$. Furthermore, from (42) and the a.e. convergence of $\rho_{\varepsilon}$ we deduce that
$\limsup_{\varepsilon\to 0}|\nabla\rho_{\varepsilon}|\leq\chi\rho\quad\mbox{a.e.\leavevmode\nobreak\ on }\{\rho>0,\,\Xi=0\}.$ (44)
In particular the sequence $(\nabla\rho_{\varepsilon}(x,t))_{\varepsilon>0}$ is bounded for a.e. $(x,t)\in\mathbb{R}^{d}\times(0,T)\cap\{\rho>0\}$ (with a bound dependent on $(x,t)$).
Given the previous two a.e. convergence relations we deduce
$\rho_{\varepsilon}^{3}\frac{\left|\nabla\rho_{\varepsilon}\right|}{\rho_{\varepsilon}}{\bf 1}_{\{\rho_{\varepsilon}>0\}}\left(\frac{\left|\nabla\rho_{\varepsilon}\right|}{\rho_{\varepsilon}}{\bf 1}_{\{\rho_{\varepsilon}>0\}}+\chi\right)\left(1-\frac{\chi\rho_{\varepsilon}}{|\nabla\rho_{\varepsilon}|}\right)_{+}{\bf 1}_{\{\rho_{\varepsilon}>0\}}\frac{\nabla\rho_{\varepsilon}}{\rho_{\varepsilon}}{\bf 1}_{\{\rho_{\varepsilon}>0\}}\rightarrow\overline{w},$
almost everywhere in $0,T]\times\mathbb{R}^{d}\setminus\{\rho=0\}$, or, which is equivalent,
${\bf 1}_{\{\rho_{\varepsilon}>0\}}\left(|\nabla\rho_{\varepsilon}|^{2}-\rho_{\varepsilon}^{2}\chi^{2}\right)_{+}\nabla\rho_{\varepsilon}\rightarrow\overline{w},$ (45)
almost everywhere in $0,T]\times\mathbb{R}^{d}\setminus\{\rho=0\}$. In fact, while there is clearly convergence a.e. on $\{\rho>0,\,\Xi>0\}$, on $\{\rho>0,\,\Xi=0\}$ the term $\nabla\rho_{\varepsilon}$ is bounded thanks to (44) while $\left(1-\frac{\chi\rho_{\varepsilon}}{|\nabla\rho_{\varepsilon}|}\right)_{+}{\bf 1}_{\{\rho_{\varepsilon}>0\}}$ tends to zero.
Notice that ${\bf 1}_{\{\rho_{\varepsilon}>0\}}$ can be eliminated from (45) since $\left(|\nabla\rho_{\varepsilon}|^{2}-\rho_{\varepsilon}^{2}\chi^{2}\right)_{+}\nabla\rho_{\varepsilon}=0$ on $\{\rho_{\varepsilon}=0\}$. Moreover, as $\rho_{\varepsilon}\rightarrow\rho$ a.e. in $(0,T)\times\mathbb{R}^{d}$ and the sequence $(\nabla\rho_{\varepsilon}(x,t))_{\varepsilon>0}$ is bounded for a.e. $(x,t)\in\mathbb{R}^{d}\times(0,T)\cap\{\rho>0\}$, we infer
@@
almost everywhere in $0,T]\times\mathbb{R}^{d}\setminus\{\rho=0\}$.
@@
$+\frac{1}{2}\int_{Q_{R,\tau,T}^{\eta}}\left[\left(|\nabla\rho_{\varepsilon}|^{2}-\rho^{2}\chi^{2}\right)_{+}-\left(|\nabla\rho|^{2}-\rho^{2}\chi^{2}\right)_{+}\right]\times\left(\lvert\nabla\rho_{\varepsilon}\rvert^{2}-\lvert\nabla\rho\rvert^{2}\right)\,{\rm d}x\,{\rm d}t\rightarrow 0.$ (46b)
Since both integrals in (46a), (46b) are nonnegative, the convergence to 0 of (46a) implies that we can find a subsequence (not relabeled) such that $\nabla\rho_{\varepsilon}\rightarrow\nabla\rho$ a.e. in $Q_{R,\tau,T}^{\eta}\setminus\{\rho\chi\geq|\nabla\rho|\}$. W.l.o.g. we can choose $\eta=1/N$, $N\in\mathbb{N}$, and assume that $Q_{R,\tau,T}^{1/N_{1}}\subset Q_{R,\tau,T}^{1/N_{2}}$ for $N_{1}<N_{2}$. Since $|Q_{R,\tau,T}\setminus Q_{R,\tau,T}^{\eta}|<\eta$, a Cantor diagonal argument yields the existence of a subsequence (not relabeled) such that $\nabla\rho_{\varepsilon}\rightarrow\nabla\rho$ a.e. in $Q_{R,\tau,T}\setminus\{\rho\chi\geq|\nabla\rho|\}$. Via another Cantor diagonal argument we conclude that (up to a subsequence)
$\nabla\rho_{\varepsilon}\rightarrow\nabla\rho\leavevmode\nobreak\ \leavevmode\nobreak\ \mbox{ a.e. in }(0,T)\times\mathbb{R}^{d}\cap\{0<\rho\chi<|\nabla\rho|\}.$ (47)
@@
$\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}\nabla\rho_{\varepsilon}\to\left(1-\frac{\chi\rho}{\lvert\nabla\rho\rvert}\right)_{+}\nabla\rho\quad\mbox{a.e.\leavevmode\nobreak\ in }(0,T)\times\mathbb{R}^{d}\cap\{0<\rho\chi<|\nabla\rho|\}.$ (48)
@@
$\hskip-17.07164pt\left[\left(|\nabla\rho_{\varepsilon}|^{2}-\rho^{2}\chi^{2}\right)_{+}-\left(|\nabla\rho|^{2}-\rho^{2}\chi^{2}\right)_{+}\right]\times\left(\lvert\nabla\rho_{\varepsilon}\rvert^{2}-\lvert\nabla\rho_{\varepsilon}\rvert^{2}\right)$
@@
$\left(|\nabla\rho_{\varepsilon}|^{2}-\rho_{\varepsilon}^{2}\chi^{2}\right)_{+}\rightarrow\left(|\nabla\rho|^{2}-\rho^{2}\chi^{2}\right)_{+}\quad\mbox{a.e. in }{(0,T)\times\mathbb{R}^{d}\cap\{\rho>0\}}.$
As a consequence we find
@@
$\mbox{a.e.\leavevmode\nobreak\ in }(0,T)\times\mathbb{R}^{d}\cap\{\rho\chi\geq|\nabla\rho|,\leavevmode\nobreak\ {\rho>0}\},$
hence we obtain
$\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}\nabla\rho_{\varepsilon}\to 0=\left(1-\frac{\chi\rho}{\lvert\nabla\rho\rvert}\right)_{+}\nabla\rho\quad\mbox{a.e.\leavevmode\nobreak\ in }(0,T)\times\mathbb{R}^{d}\cap\{\rho\chi\geq|\nabla\rho|,\leavevmode\nobreak\ {\rho>0}\}.$ (49)
To identify the a.e. limit on the set $(0,T)\times\mathbb{R}^{d}\cap\{\rho=0\}$ we proceed as follows. Consider
$\left\|\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}\nabla\rho_{\varepsilon}\,{\bf 1}_{\{\rho=0\}}\right\|_{4/3}\leq 2\|\nabla\sqrt{\rho_{\varepsilon}}\|_{2}\|\sqrt{\rho_{\varepsilon}}{\bf 1}_{\{\rho=0\}}\|_{4}=2\|\nabla\sqrt{\rho_{\varepsilon}}\|_{2}\|\rho_{\varepsilon}{\bf 1}_{\{\rho=0\}}\|_{2}^{1/2}.$
Given that $\|\nabla\sqrt{\rho_{\varepsilon}}\|_{2}$ is bounded and $\rho_{\varepsilon}{\bf 1}_{\{\rho=0\}}\to\rho{\bf 1}_{\{\rho=0\}}=0$ strongly in $L^{2}(\mathbb{R}^{d}\times(0,T))$, we conclude that $\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}\nabla\rho_{\varepsilon}\to 0$ strongly in $L^{4/3}(\mathbb{R}^{d}\times(0,T)\cap\{\rho=0\})$. In particular, up to subsequences
$\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}\nabla\rho_{\varepsilon}\rightarrow 0=\left(1-\frac{\chi\rho}{\lvert\nabla\rho\rvert}\right)_{+}\nabla\rho\quad\mbox{a.e.\leavevmode\nobreak\ in }(0,T)\times\mathbb{R}^{d}\cap\{\rho=0\}.$ (50)
Putting (48)–(50) together yields the desired conclusion
@@
and we are finally able to identify the weak $L^{2}$ limit of the term $\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}\nabla\rho_{\varepsilon}$:
$\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}\nabla\rho_{\varepsilon}\rightharpoonup\left(1-\frac{\chi\rho}{\lvert\nabla\rho\rvert}\right)_{+}\nabla\rho\quad\mbox{weakly in }L^{2}((0,T)\times\mathbb{R}^{d}).$ (51)
4.3. The vanishing viscosity limit
We have now all the ingredients to pass to the limit in (19). From (22) we obtain
$\|\partial_{t}\rho_{\varepsilon}\|_{L^{2}(0,T;H^{-1}(\mathbb{R}^{d}))}\leq C\|\rho_{{\rm in},\varepsilon}\|_{L^{2}(\mathbb{R}^{d})}.$
Since $\rho_{{\rm in},\varepsilon}$ is uniformly bounded in $L^{2}(0,T;\mathbb{R}^{d})$, the last inequality yields
@@ 4.3. The vanishing viscosity limit
Convergence (51) implies that
$\int\int\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}\nabla\rho_{\varepsilon}\nabla\varphi\;dxdt\to\int\int\left(1-\frac{\chi\rho}{\lvert\nabla\rho\rvert}\right)_{+}\nabla\rho\nabla\varphi\;dxdt.$
The viscous terms vanish, since
$\lim_{\varepsilon\rightarrow 0}\left|\varepsilon\int_{0}^{T}\int_{\mathbb{R}^{d}}\nabla\rho_{\varepsilon}\nabla\varphi\,{\rm d}x\,{\rm d}t+\varepsilon\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\varphi\,{\rm d}x\,{\rm d}t\right|$ $\leq\lim_{\varepsilon\rightarrow 0}\varepsilon C\|\rho_{{\rm in},\varepsilon}\|_{L^{2}(0,T;\mathbb{R}^{d})}\|\varphi\|_{L^{2}(0,T;H^{1}(\mathbb{R}^{d}))}$
@@
Hence, we have constructed a weak solution $\rho$ to (11). In the next two subsection we show that such solution is unique and converges strongly in $L^{p}$ to the initial data $\rho_{in}$.
4.4. Obtaining the initial data
@@
$\left\{u\in L^{2}(0,T;H^{1}(\mathbb{R}^{d}))\leavevmode\nobreak\ \leavevmode\nobreak\ :\leavevmode\nobreak\ \leavevmode\nobreak\ \partial_{t}u\in L^{2}(0,T;H^{-1}(\mathbb{R}^{d}))\right\},$
@@
$\lim_{\varepsilon\to 0}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}(t)\varphi dx=\int_{\mathbb{R}^{d}}\rho(t)\varphi dx\quad\forall\varphi\in H^{1}(\mathbb{R}^{d}).$ (52)
Also remember that the initial condition for the approximated problem (19) is satisfied in the sense of (21).
Define for brevity $A_{\varepsilon}:=\varepsilon+\left(1-\frac{\chi\rho_{\varepsilon}}{|\nabla\rho_{\varepsilon}|}\right)_{+}$, $A:=\left(1-\frac{\chi\rho}{|\nabla\rho|}\right)_{+}$. Choose $\varphi\in H^{1}(\mathbb{R}^{d})$ independent of time in (19). It follows
@@
$F_{M}(s)=\int_{0}^{s}f_{M}(\sigma)d\sigma,\quad f_{M}(s)=\begin{cases}ps^{p-1}&0\leq s\leq M\\ p(p-1)M^{p-2}(s-M)+pM^{p-1}&s>M\end{cases}.$
$\int_{\mathbb{R}^{d}}F_{M}(\rho(t_{2}))dx-\int_{\mathbb{R}^{d}}F_{M}(\rho(t_{1}))dx$ $=\int_{t_{1}}^{t_{2}}\int_{\mathbb{R}^{d}}Af_{M}^{\prime}(\rho)|\nabla\rho|^{2}dxdt$
$=\frac{4}{p^{2}}\int_{t_{1}}^{t_{2}}\int_{\mathbb{R}^{d}}Af_{M}^{\prime}(\rho)\rho^{2-p}|\nabla\rho^{p/2}|^{2}dxdt.$
$\int_{\mathbb{R}^{d}}\rho(t_{2})^{p}dx-\int_{\mathbb{R}^{d}}\rho(t_{1})^{p}dx$ $=\frac{4(p-1)}{p}\int_{t_{1}}^{t_{2}}\int_{\mathbb{R}^{d}}A|\nabla\rho^{p/2}|^{2}dxdt,$
for a.e. $t_{1},t_{2}\in[0,T]$. It follows that the function $t\mapsto\|\rho(t)\|_{L^{p}(\mathbb{R}^{d})}^{p}$ belongs to $W^{1,1}(0,T)$, hence up to changing the values of $\rho:[0,T]\to L^{p}(\mathbb{R}^{d})$ in a zero measure set it holds that $t\mapsto\|\rho(t)\|_{L^{p}(\mathbb{R}^{d})}^{p}$ is continuous in $[0,T]$. As a consequence the bound $\|\rho(t)\|_{L^{p}(\mathbb{R}^{d})}\leq\|\rho_{in}\|_{L^{p}(\mathbb{R}^{d})}$ holds for every $t\in[0,T]$. In particular $\rho(t)$ is weakly convergent (up to subsequences) in $L^{p}(\mathbb{R}^{d})$ as $t\to 0$; however, its weak limit has to be $\rho_{\in}$ due to (53), hence $\rho(t)\rightharpoonup\rho_{in}$ weakly in $L^{p}(\mathbb{R}^{d})$ as $t\to 0$ (no longer up to subsequences). It follows
$\|\rho_{in}\|_{L^{p}(\mathbb{R}^{d})}\leq\liminf_{t\to 0}\|\rho(t)\|_{L^{p}(\mathbb{R}^{d})}\leq\limsup_{t\to 0}\|\rho(t)\|_{L^{p}(\mathbb{R}^{d})}\leq\|\rho_{in}\|_{L^{p}(\mathbb{R}^{d})},$
implying that $\|\rho(t)\|_{L^{p}(\mathbb{R}^{d})}\to\|\rho_{in}\|_{L^{p}(\mathbb{R}^{d})}$ as $t\to 0$. This fact and the weak convergence $\rho(t)\rightharpoonup\rho_{in}$ in $L^{p}(\mathbb{R}^{d})$, imply that $\rho(t)\rightarrow\rho_{{\rm in}}$ strongly in $L^{p}(\mathbb{R}^{d})$ as $t\to 0$.
4.5. Uniqueness
Let $\rho_{1}$ and $\rho_{2}$ be two solutions to (11) with the same initial data $\rho_{in}$. As in Section 4.1, we subtract equation (11) for $\rho_{2}$ tested with $\varphi_{2}=\frac{\rho_{1}+\sigma}{\rho_{2}+\sigma}{-1}=\frac{\rho_{1}-\rho_{2}}{\rho_{2}+\sigma}$ to the equation for $\rho_{1}$ tested with $\varphi_{1}=\log\left(\frac{\rho_{1}+\sigma}{\rho_{2}+\sigma}\right)$. Following the same computations as in (26), we get
$=\int_{0}^{T}\frac{d}{dt}\int_{\mathbb{R}^{d}}\left({\rho_{1}+\sigma}\right)\log{\left(\frac{\rho_{1}+\sigma}{\rho_{2}+\sigma}\right)}-\rho_{1}+\rho_{2}\,{\rm d}x\,{\rm d}t=\int_{0}^{T}\frac{d}{dt}H[\rho_{1}+\sigma|\rho_{2}+\sigma]\;dt.$
On the other hand, substituting $\partial_{t}\rho_{1}$ and $\partial_{t}\rho_{2}$ with the respective right hand side of (11), after some integration by parts we obtain:
$\int_{0}^{T}\int_{\mathbb{R}^{d}}\left[\varphi_{1}\partial_{t}\rho_{1}-\varphi_{2}\partial_{t}\rho_{2}\right]\,{\rm d}x\;dt$ (54)